Solve the linear system:
5x-y=7
x=-4y+16

Answer by Alan3354(69443)

5x-y=7
x=-4y+16
---------
Sub for x in the 1st eqn
5(-4y+16) - y = 7
-20y + 80 - y = 7
-21y = -73
y = 73/21
--------------
Sub for y in the 2nd eqn